CAMILLEANN:

A charge of 5 x10-9 C is attracted by a charge of –3x10-7 C with a force of 0.135 N. How far apart are they? (Hint: manipulate the equation to get an expression for d)

Tranquility:

You have to use Coulomb's law

1 attachment
Tranquility:

\(\Large F = k\dfrac{Q_1 Q_2}{d^2}\) F = Force = 0.135 N charge 1 = Q1 = 5 * 10^-9 C charge 2 = Q2 = -3 * 10^-7 C k = Coulomb's constant = 8.9875 * 10^9 N*m^2/C^2 distance = d = ?? Solve for d which is the distance between those two charges